OpenStudy (anonymous):

use the implicit differentiation to find the equation of the tangent line to the curve xy^3+xy=10at point (5,1)

OpenStudy (amistre64):

what do you know about implicit differentation?

OpenStudy (anonymous):

differentiate it and subs the point (5,1) . am i right? @amistre

OpenStudy (amistre64):

yep, thats right think

OpenStudy (amistre64):

differentiate like normal, just keep your x' and y' in the problem...

OpenStudy (anonymous):

y-1=-(x-5)/10

OpenStudy (amistre64):

D(xy^3) + D(xy) = D(10) product rule ... product rule....constant rule

OpenStudy (anonymous):

simplify and differentiate, 10y-10=-x-5 => x+10y=5now try to differentiate...will that do @amistre?

OpenStudy (amistre64):

you CAN try to seperate it all, but its really just easier to work the problam as it...

OpenStudy (amistre64):

(x'y + xy') + (x'y + xy') = D(10) = 0

OpenStudy (amistre64):

x'(y^3) + x(3y^2)y' + x'y + xy' = 0

OpenStudy (anonymous):

yeah this is fine even..whatever u find easy diddy :)

OpenStudy (amistre64):

x' = dx/dx = 1 like any good fraction... now solve for y'

OpenStudy (anonymous):

hey catch u later ..gotta go :) bye all

OpenStudy (amistre64):

Ciao think :)

OpenStudy (anonymous):

thanks @ thinker

OpenStudy (anonymous):

ur welcome ..@diddy, all the best and ciao amistre...ttyl maybe after 2 days :)

OpenStudy (amistre64):

Its important to realize that there is nothing that special when doing implicits.. for example: D(x^4) = x' 4x^3 D(3y^2) = y' 6y D(3xy) = 3xy' + x'3y D(y) = y' D(x) = x' its all normal rules

OpenStudy (anonymous):

oki so i used this form amistre x'(y^3) + x(3y^2)y' + x'y + xy' = 0 and i came up with 2y^3/3xy^2+1, but i'm not sure that's right

OpenStudy (amistre64):

Let me double check the work... x'(y^3) + x(3y^2)y' + x'y + xy' = 0 y^3 + y'(3xy^2) + y + y'x = 0 y'(3xy^2) + y'x = -y^3 - y y'(3xy^2 + x) = -y^3 -y -y^3 -y y' = ----------- 3xy^2 + x did I make any mistakes in this?

OpenStudy (anonymous):

You didn't :)

OpenStudy (amistre64):

yay!!... lately, thats an accomplishment :)

OpenStudy (anonymous):

yea i think that's right, make sense now thanks :)

OpenStudy (amistre64):

now plug in your x and y values to determine the slope of the tangent line at that point :)

OpenStudy (anonymous):

ook got it

OpenStudy (amistre64):

good job :)

OpenStudy (anonymous):

YAY i got it!!! thanks u're the best :)

OpenStudy (amistre64):

now, now... youre only saying that cause its true lol :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i'm having some trouble with this one... Find the slope of the tangent line to the curve -1x^2-2xy-3y^3=-185 at the point (-1,4).

OpenStudy (anonymous):

never mind, got it!!!